OpenStudy (anonymous):

The price demand equation and cost functions for a new product are x=9000-30p and c(x)=30x+150000? 0 < equal x < equal 90000 express the price p as a function of the demand x. answer is P=300-x/30 someone explain me how they got 300?

OpenStudy (zpupster):

solve the first equation for P x=9000-30P x-9000=-30P x-9000/-30 =-30P/-30 x/-30 +(-9000/-30) = P rearranging P=300-x/30

OpenStudy (anonymous):

thank u ;)